OpenStudy (anonymous):

alan drives his shiny, new car around a racertack everyday. the race track has a diameter of 60 meters/ one day he manages to go all the way around the track in 12pi seconds. what is his average speed in meters per second?

OpenStudy (anonymous):

circumference of the race track is required. is the race track a circle?

OpenStudy (anonymous):

it doesn't say but I would think yes because it gives diameter

OpenStudy (anonymous):

I also think so so whats the circumference of a circle ?

OpenStudy (anonymous):

2pir

OpenStudy (anonymous):

so the meters he makes when going one round is round = pi * 60m

OpenStudy (anonymous):

he makes the round in 12pi seconds, so: \[v = \frac{ \pi \times 60m }{ 12\pi s}\]

OpenStudy (anonymous):

so do i leave it like that or simplify?

OpenStudy (anonymous):

\[\frac{ 5m }{ s} \] ?

OpenStudy (anonymous):

yes I would get v= 5 m/s

OpenStudy (anonymous):

ok thanks